OpenStudy (anonymous):

write the equation in standard form for the ellipse: 16x^2+48x+4y^2-20y+57=0 this is the answer that i got 4(x+3/2)^2 + (y-5/2)^2 = 1 im not sure if the 4 is suppose to be there

OpenStudy (anonymous):

It's correct

OpenStudy (anonymous):

but doesn't the denominators needs to be different?

OpenStudy (anonymous):

4(x+3/2)^2 + (y-5/2)^2 = 1 => 4(2x+3)^2 + (2y-5)^2 = 4

OpenStudy (anonymous):

your answer was correct , but there are some modification ,just take LCM through out and take the denominator to the right hand side

OpenStudy (anonymous):

for an ellipse the right hand side must be 1

OpenStudy (anonymous):

if you don't like the 4 then move it to the denominator as (1/2)^2

OpenStudy (anonymous):

(x+3/2)^2/.5^2 + (y-5/2)^2 = 1 and technically the denominators don't have to be different, because a circle is also an ellipse